# Allocation Multiplicity: Evaluating the Promises of the Rashomon Set

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: noneccs: Computing methodologies Philosophical/theoretical foundations of artificial intelligence
@@ 1. Introduction
Many decision-making settings involve allocating a fixed number of resources or opportunities among individuals. A company might have hundreds of job applicants but only ten job openings to fill, or a healthcare program may be able to enroll only 100 patients out of thousands. In these cases, there are often more qualified candidates than resources or goods available. This creates allocation multiplicity: the existence of many equal-utility selections of individuals. For example, there are over 3.2 million possible ways to choose 10 qualified people from 25 qualified applicants. The number of possible equal-utility allocations increases even more if we assume that decision-makers will make mistakes and choose some unqualified individuals.
@@
Figure 1. Model Multiplicity vs Allocation Multiplicity. Suppose that 10 individuals with different qualifications compete for 5 scarce resources. 2 models from the Rashomon set yield different predictions or risk scores, leading to 2 different equal-utility allocations in which 4 out of 5 selected individuals are qualified. In this setting, there are 60 such equal-utility allocations based on (${6\choose 4}\cdot{4\choose 1}$.
@@
Figure 2. Rashomon Set of Models vs Equal-Utility Allocations. The Rashomon set of all equally-good models is not enumerable, but empirical methods can sample from it. This sample of Rashomon models will lead to some recovered allocations, which may not faithfully represent the full space of equal-utility allocations due to (1) train-test differences, (2) deterministic selection of the lowest risk individuals, and (3) structural biases that cause all near-optimal models to view some qualified individuals as inherently risky.
@@
In this work, we argue that the Rashomon set primarily achieves these aims to the extent that it supports allocation multiplicity. Figure 1 shows that model multiplicity results in different individual-level predictions, which lead to different equal-utility allocations. However, the Rashomon set of models will not yield the full space of equal-utility allocations for several reasons. First, it is unfeasible to enumerate all possible models in the Rashomon set for unbounded hypothesis classes, so we must rely on empirical methods to sample from it (Li et al., 2024; Semenova et al., 2022). This sample of Rashomon models is typically characterized by similar training or validation performance (Watson-Daniels et al., 2023a; Marx et al., 2020), which differs from deployment utility in a resource-constrained setting. Second, decision-makers usually allocate resources by deterministically selecting individuals with the lowest risk, overlooking stochastic mappings between predictions and decisions (Jain et al., 2024a). Third, structural biases in the measurable features and available training data may cause all models in the Rashomon set to consider some qualified individuals as inherently risky. As Figure 2 illustrates, these reasons may lead to recovered allocations from a sample of Rashomon models that are highly unrepresentative of the full space of equal-utility allocations.
@@
The concept of allocation multiplicity enables us to evaluate how well the Rashomon set fulfills its promises of less discriminatory algorithms, reduced outcome homogenization, and fairer decisions through model ensembles or reconciliation. Through a case study of healthcare allocations, we explore various methods for sampling from the Rashomon set and compare their recovered allocations to the full space of equal-utility allocations. Our results demonstrate that: (1) many less discriminatory allocations are not found, (2) outcome homogenization persists in recovered allocations, and (3) the allocation from an ensemble model is unrepresentative of the full space of equal-utility allocations. These findings underscore that allocation multiplicity should serve as a new benchmark for model multiplicity and that recovering equal-utility allocations is essential to fulfilling the promises of the Rashomon set.
@@ 2. Model Multiplicity and the Rashomon Set
A growing literature discusses the phenomenon of model multiplicity: the existence of many equally-good models that differ in their individual-level predictions or other properties (Black et al., 2022; Marx et al., 2020; Watson-Daniels et al., 2023b). Equally-good models are models with similar performance for a chosen metric (e.g. accuracy or loss) on a fixed dataset (Ganesh, 2024; Watson-Daniels et al., 2023a). Model multiplicity can arise due to procedural multiplicity, which refers to different choices along the model-building pipeline (Black et al., 2022, 2023). This includes decisions about the problem formulation, feature selection, training data, and model class (Watson-Daniels et al., 2023a; Meyer et al., 2023). However, even when these choices are fixed, model multiplicity can still arise due to uncertainty or noise in the data as well as randomness in the model training process. For example, the data may have been generated through noisy or nondeterministic processes, such as human decision-making (Rudin et al., 2024; Semenova et al., 2023). The data may contain insufficient information for a model to learn generalizable domain knowledge, a problem known as underspecification (D’Amour et al., 2022). Additionally, randomness in the model training process, such as weight initialization, and random seeds, further contribute to model multiplicity (Ganesh et al., 2023).
The Rashomon set [^1] Breiman (2001) named the Rashomon set after the movie “Rashomon,” directed by Akira Kurosawa, in which four different perspectives on the same murder fail to converge to a single truth. consists of all models with near-optimal performance (Breiman, 2001; Fisher et al., 2019; Rudin et al., 2024). While this set generally includes all “good” models, the set can also be defined for a specific hypothesis class, evaluation dataset, and loss function (Hsu and Calmon, 2022; Semenova et al., 2022; Watson-Daniels et al., 2023a). This allows for analyses of its size through metrics like the Rashomon ratio (Semenova et al., 2022) and Rashomon capacity (Hsu and Calmon, 2022). Several empirical methods exist to iterate over the Rashomon set for simple hypothesis classes or sample from it for more complex model types. For example, the Rashomon set is enumerable for decision-trees with a low regularized risk value (Xin et al., 2022), sparse generalized additive models (Zhong et al., 2023), and sparse scoring systems (Liu et al., 2022). Complex hypothesis classes like neural networks have methods to explicitly search for Rashomon models, such as adversarial weight perturbation (Hsu and Calmon, 2022) and dropout-based exploration (Hsu et al., 2024). Other approaches may implicitly produce Rashomon models, such as training on bootstrap datasets (Cooper et al., 2024) or changing random seeds (Ganesh et al., 2023).
@@ 2.1. Promises of the Rashomon Set
Model multiplicity implies that for a given model and a fairness metric, there may exist “less discriminatory algorithms” (LDAs) that are equally-accurate (Black et al., 2022, 2024b). Several works have empirically shown that LDAs exist for fixed populations (Auerbach et al., 2024; Gillis et al., 2024; Coston et al., 2021), but the chosen LDA may be more discriminatory in unseen deployment settings (Black et al., 2024a; Laufer et al., 2024). The goal of searching for LDAs is usually to find a model that improves on group fairness metrics like statistical parity or disparities in error rates without compromising accuracy. This motivates using the Rashomon set to find LDAs (Black et al., 2022; Rudin et al., 2024). In domains such as credit, employment, and housing, decision-makers may have a duty to search for LDAs based on U.S. anti-discrimination laws (Black et al., 2024b). Under the disparate impact doctrine, defendants can justify using a model with discriminatory effects by claiming the model serves a “business necessity,” which often relates to its performance. But even with this justification, defendants may still face liability if the plaintiffs can show there exists a “less discriminatory alternative” that serves the same purpose. In some cases, the courts have placed the burden on defendants to conduct a “reasonable search” for LDAs (Black et al., 2024b; Laufer et al., 2024).
@@ 3.1. Preliminaries and Problem Formulation
Risk-averse decision-makers aim to maximize utility by selecting individuals with the highest probability of being qualified. While the true probability $p_{i}=\mathbb{P}(o^{*}_{i}=1)$ is unknowable [^5] Individual probabilities misrepresent qualifications because the events in question are typically realized only once (Dawid, 2017; Dwork et al., 2021; Roth et al., 2023)., the conditional probability $p(x_{i})=\mathbb{P}(o^{*}_{i}=1\,|\,x_{i})$ can be estimated using data. The validity of taking $p(x_{i})$ to be an estimate of $p_{i}$ depends on the choice of features and how well qualifications can be predicted from them. Setting these specifications aside, machine learning models can estimate $p(x_{i})$ based on features $x_{i}$ from a given domain $\mathcal{X}$.
@@
A mapping =$\mathcal{M}=\{(\hat{p}(x_{1}),o_{1}),(\hat{p}(x_{2}),o_{2}),\dots,(\hat{p}(x_{n}),o_{n})\}$ characterizes the relationship between a model $\mathbf{\hat{p}}$ and allocation $\mathbf{o}$. In a top $k$ mapping, $o_{i}=1$ if and only if $\,\hat{p}(x_{i})$ is in the $k$ highest predictions.
@@ 3.2. From Model Multiplicity to Allocation Multiplicity
$\mathcal{R}_{\epsilon}(\mathbf{\hat{p}_{0}}):=\{\mathbf{\hat{p}}\in\mathcal{H}:\mathbb{L}(\mathbf{\hat{p}};\mathcal{D_{\texttt{train}}})\leq\mathbb{L}(\mathbf{\hat{p}_{0}};\mathcal{D_{\texttt{train}}})+\epsilon\}$
@@ 3.3. Quantifying Allocation Multiplicity
[…]
@@
[…]
@@ 4. Why Does Allocation Multiplicity Matter?
Now that we have seen how many equal-utility allocations are possible, we discuss why allocation multiplicity matters. First, risk-averse decision-makers might assume that all or most good equal-utility allocations will be produced by the set of equally-good models. Allocation multiplicity highlights that “riskier” allocations can be chosen without sacrificing utility. Second, decision-makers who prioritize normative ideals other than utility may value allocation multiplicity for reasons aligned with their principles. Third, those who view the Rashomon set positively for its potential benefits may find that the space of all equal-utility allocations is what actually delivers those benefits.
4.1. From the Perspective of Risk-Aversion
@@
Risk-averse decision-makers will usually choose the allocation with the highest expected utility based on a model’s predictions. And from the perspective of the information they have at the time, this might be the most utility-maximizing step they can take. But model multiplicity means that even in the set of models that are built with the same problem formulation and trained on the same data, there are many models that could minimize risk to an equivalent extent, thereby delivering equivalent expected utility to the risk-averse. We take the next step, pointing out that realized utility, not the expected utility, is what ultimately matters to most decision-makers. While our current information does not always allow us to find all the equal-utility allocations, the knowledge that they exist should spur decision-makers to collect the additional information needed to recover more, and a more even sampling, of the equal-utility allocations. Comparing each highest-expected-utility model with the set of all possible equal-utility allocations in the training data is a tool to highlight the shortcomings of the current training data and the inherent limitations of allocating based on solely on minimizing risk.
From the Perspective of Claims
Allocation multiplicity has a positive role in ensuring that the use of machine learning models to make decisions lives up to our normative commitments. For example, we might be morally committed to fulfilling claims that individuals have to goods or outcomes for reasons of fairness (Broome, 1990). According to the claims view, an individual has a claim to an organ if she needs it more than others; a claim to a job if she merits it; or a claim to a good if she was promised it and therefore deserves it (Broome, 1990). If resources are scarce, not all claims can be fulfilled. Our comparison of model multiplicity to allocation multiplicity show that drawing predictions only from the Rashomon set models over-respects the claims of those who appear less risky and under-respects the claims of those who appear riskier. But those with equal claims deserve equal chances at the good. Although we may not always have the information needed to predict allocations correctly, the claims framework suggests that we should (1) seek out additional information that would allow us to better respect all claims and (2) be prepared to give “compensatory” goods to those whose claims we failed to respect.
From the Perspective of the Worst-Off
Another variant of utilitarianism that can guide us in creating and choosing allocations is prioritarianism. Prioritarianism holds that when allocating resources, we should prioritize improvements to the well-being of the worst-off (Arneson, 2022). Considering equal-utility allocations allows prioritiarians to help the worst-off in two ways. First, within the set of apparently equal-utility allocations, there are likely allocations that give more resources to those who suffer the most. Thus just as the search for the LDA can be used to increase fairness by choosing the fairest of the equally-good models, the prioritarian may search the equal-utility allocation space for an allocation that prioritizes the worst-off. Second, the shift in perspective from model multiplicity to allocation multiplicity suggests that many of the “multiple” models have allocations that are concentrated on the least-risky individuals. Those individuals are likely to be the best-off, not the worst off. So prioritarianism would suggest using the allocation-multiplicity perspective to spur the search for the seemingly-risky people who would be most benefited by an allocation.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732138isbn: 979-8-4007-1482-5/2025/06ccs: Computing methodologies Philosophical/theoretical foundations of artificial intelligence
@@ 1. Introduction
Many decision-making settings involve allocating a fixed number of resources or opportunities among individuals. A company might have hundreds of job applicants but only ten job openings to fill; a healthcare program may be able to enroll only a 100 patients out of thousands. In these cases, there are often more qualified candidates than resources or goods available. This creates allocation multiplicity: the existence of many equal-utility selections of individuals. For example, there are over 3.2 million possible ways to choose 10 qualified people from 25 qualified applicants. The number of possible equal-utility allocations increases even more if we assume that decision-makers will make mistakes and choose some unqualified individuals.
@@
Figure 1. Suppose that 10 individuals with different qualifications compete for 5 scarce resources. 2 models from the Rashomon set yield different predictions or risk scores, leading to 2 different equal-utility allocations in which 4 out of 5 selected individuals are qualified. In this setting, there are 60 such equal-utility allocations based on ${6\choose 4}\cdot{4\choose 1}$.
[…]
Suppose that 10 individuals with different qualifications compete for 5 scarce resources. 2 models from the Rashomon set yield different predictions or risk scores, leading to 2 different equal-utility allocations in which 4 out of 5 selected individuals are qualified. In this setting, there are 60 such equal-utility allocations based on 6 choose 4 cdot 4 choose 1.
@@
Figure 2. The Rashomon set of all equally-good models is not enumerable, but empirical methods can sample from it. This sample of Rashomon models will lead to recovered allocations that may not faithfully represent the full space of equal-utility allocations due to (1) train-test data differences, (2) deterministic selection of the lowest risk individuals, and (3) structural biases that cause all near-optimal models to view some qualified individuals as inherently risky.
[…]
The Rashomon set of all equally-good models is not enumerable, but empirical methods can sample from it. This sample of Rashomon models will lead to recovered allocations that may not faithfully represent the full space of equal-utility allocations due to (1) train-test data differences, (2) deterministic selection of the lowest risk individuals, and (3) structural biases that cause all near-optimal models to view some qualified individuals as inherently risky.
@@
In this work, we argue that the Rashomon set primarily achieves these aims to the extent that it supports allocation multiplicity. Figure 1 shows that model multiplicity results in different individual-level predictions, which lead to different equal-utility allocations. However, the Rashomon set of models will not yield the full space of equal-utility allocations for several reasons. First, it is infeasible to enumerate all possible models in the Rashomon set for unbounded hypothesis classes, so we must rely on empirical methods to sample from it (Li et al., 2024; Semenova et al., 2022). This sample of Rashomon models is typically characterized by similar training or validation performance (Watson-Daniels et al., 2023a; Marx et al., 2020), which differs from deployment utility in a resource-constrained setting. Second, decision-makers usually allocate resources by deterministically selecting individuals with the lowest risk, overlooking stochastic mappings between predictions and decisions (Jain et al., 2024a). Third, structural biases in the measurable features and available training data may cause all models in the Rashomon set to consider some qualified individuals as inherently risky. As Figure 2 illustrates, these reasons may lead to recovered allocations from a sample of Rashomon models that are highly unrepresentative of the full space of equal-utility allocations.
@@
The concept of allocation multiplicity enables us to evaluate how well the Rashomon set fulfills its promises of less discriminatory algorithms, reduced outcome homogenization, and fairer decisions through model ensembles or reconciliation. Through a case study of healthcare allocations, we explore various methods for sampling from the Rashomon set and compare their recovered allocations to the full space of equal-utility allocations. Our results demonstrate that: (1) many less discriminatory allocations are not found, (2) outcome homogenization persists in recovered allocations, and (3) ensemble models are unrepresentative of the full space of equal-utility allocations. These findings underscore that recovering diverse equal-utility allocations is essential to fulfilling the promises of the Rashomon set.
@@ 2. Model Multiplicity and the Rashomon Set
A growing literature discusses the phenomenon of model multiplicity: the existence of many equally-good models that differ in their individual-level predictions or other properties (Black et al., 2022; Marx et al., 2020; Watson-Daniels et al., 2023b). Equally-good models are models with similar performance for a chosen metric (e.g. accuracy or loss) on a fixed dataset (Ganesh, 2024; Watson-Daniels et al., 2023a). Model multiplicity can arise due to procedural multiplicity, which refers to different choices along the model-building pipeline (Black et al., 2022, 2023). This includes decisions about the problem formulation, feature selection, training data, and model class (Watson-Daniels et al., 2023a; Meyer et al., 2023). However, even when these choices are fixed, model multiplicity can still arise due to uncertainty or noise in the data as well as randomness in the model training process. For example, the data may have been generated through noisy or nondeterministic processes, such as human decision-making (Rudin et al., 2024; Semenova et al., 2023). The data may contain insufficient information for a model to learn generalizable domain knowledge, a problem known as underspecification (D’Amour et al., 2022). Additionally, randomness in the model training process, such as weight initialization and random seeds, further contribute to model multiplicity (Ganesh et al., 2023).
The Rashomon set [^1] Leo Breiman (2001) named the Rashomon set after the movie “Rashomon,” directed by Akira Kurosawa, in which four different perspectives on a murder fail to converge on a single truth. consists of all models with near-optimal performance (Breiman, 2001; Fisher et al., 2019; Rudin et al., 2024). While this set generally includes all “good” models, the set can also be defined for a specific hypothesis class, evaluation dataset, and loss function (Hsu and Calmon, 2022; Semenova et al., 2022; Watson-Daniels et al., 2023a). This allows for analyses of its size through metrics like the Rashomon ratio (Semenova et al., 2022) and Rashomon capacity (Hsu and Calmon, 2022). Several empirical methods exist to iterate over the Rashomon set for simple hypothesis classes or sample from it for more complex model types. For example, the Rashomon set is enumerable for decision-trees with a low regularized risk value (Xin et al., 2022), sparse generalized additive models (Zhong et al., 2023), and sparse scoring systems (Liu et al., 2022). Complex hypothesis classes like neural networks have methods to explicitly search for Rashomon models, such as adversarial weight perturbation (Hsu and Calmon, 2022) and dropout-based exploration (Hsu et al., 2024). Other approaches may implicitly produce Rashomon models, such as training on bootstrap datasets (Cooper et al., 2024) or changing random seeds (Ganesh et al., 2023).
@@ 2.1. Promises of the Rashomon Set
Model multiplicity implies that for a given model and a fairness metric, there may exist “less discriminatory algorithms” (LDAs) that are equally-accurate (Black et al., 2022, 2024b) but less discriminatory according to the fairness metric. Several works have shown empirically that LDAs exist for fixed populations (Auerbach et al., 2024; Gillis et al., 2024; Coston et al., 2021), but the chosen LDA may be more discriminatory in unseen deployment settings (Black et al., 2024a; Laufer et al., 2024). The goal of searching for LDAs is usually to find a model that improves on group fairness metrics like statistical parity or disparities in error rates without compromising accuracy. This motivates using the Rashomon set to find LDAs (Black et al., 2022; Rudin et al., 2024). In domains such as credit, employment, and housing, decision-makers may have a duty to search for LDAs based on U.S. anti-discrimination laws (Black et al., 2024b). Under the disparate impact doctrine, defendants can justify using a model with discriminatory effects by claiming the model serves a “business necessity,” which often relates to its performance. But even with this justification, defendants may still face liability if the plaintiffs can show there exists a “less discriminatory alternative” that serves the same purpose. In some cases, the courts have placed the burden on defendants to conduct a “reasonable search” for LDAs (Black et al., 2024b; Laufer et al., 2024).
@@ 3.1. Preliminaries and Problem Formulation
Risk-averse decision-makers aim to maximize utility by selecting individuals with the highest probability of being qualified. While the true probability $p_{i}=\mathbb{P}(o^{*}_{i}=1)$ is unknowable [^5] Individual probabilities misrepresent qualifications because the events in question are typically realized only once (Dawid, 2017; Dwork et al., 2021; Roth et al., 2023)., the conditional probability $p(x_{i})=\mathbb{P}(o^{*}_{i}=1\,|\,x_{i})$ can be estimated using data. The validity of taking $p(x_{i})$ to be an estimate of $p_{i}$ depends on the choice of features and how well qualifications can be predicted from them. Setting these specifications aside, machine learning models [^6] Our focus is on non-causal predictions of individual risk or utility because many decision-makers allocate resources in this way (Barocas et al., 2023). However, allocation multiplicity may still arise in a causal setting, if many individuals have similar treatment effects. can estimate $p(x_{i})$ based on features $x_{i}$ from a given domain $\mathcal{X}$.
@@
A mapping $\mathcal{M}=\{(\hat{p}(x_{1}),o_{1}),(\hat{p}(x_{2}),o_{2}),\,\dots$
$(\hat{p}(x_{n}),o_{n})\}$ characterizes the relationship between a model $\mathbf{\hat{p}}$ and allocation $\mathbf{o}$. In a top $k$ mapping, $o_{i}=1$ if and only if $\hat{p}(x_{i})$ is in the $k$ highest predictions.
@@ 3.2. From Model Multiplicity to Allocation Multiplicity
$\mathcal{R}_{\epsilon}(\mathbf{\hat{p}_{0}}):=\{\mathbf{\hat{p}}\in\mathcal{H}:L(\mathbf{\hat{p}};\mathcal{D_{\texttt{train}}})\leq L(\mathbf{\hat{p}_{0}};\mathcal{D_{\texttt{train}}})+\epsilon\}$
@@ 3.3. Quantifying Allocation Multiplicity
0.85 0.50 $7\text{\times}{10}^{11}$ $2\text{\times}{10}^{19}$ $3\text{\times}{10}^{17}$ $9\text{\times}{10}^{125}$ $6\text{\times}{10}^{203}$ $9\text{\times}{10}^{180}$
0.95 0.50 $\text{\times}{10}^{11} 1×1$$\text{\times}{10}^{17}$ $4\text{\times}{10}^{08}$ $4\text{\times}{10}^{115}$ $\text{\times}{10}^{174} 1×1$$\text{\times}{10}^{84}$
0.85 0.75 $5\text{\times}{10}^{12}$ $3\text{\times}{10}^{22}$ $2\text{\times}{10}^{27}$ $3\text{\times}{10}^{137}$ $4\text{\times}{10}^{237}$ $3\text{\times}{10}^{286}$
0.95 0.75 $3\text{\times}{10}^{12}$ $4\text{\times}{10}^{21}$ $7\text{\times}{10}^{23}$ $2\text{\times}{10}^{132}$ $8\text{\times}{10}^{222}$ $6\text{\times}{10}^{246}$
@@
Table 1 computes the number of equal-utility allocations for a baseline allocation with varying utility $k^{\prime}/k$, qualification rate $n^{\prime}/n$, and selection rate $k/n$. Even when there are only 100 individuals, the number of equal-utility allocations is exponentially high. For example, there are $2\text{\times}{10}^{19}$ equal-utility allocations with utility 0.85, qualification rate 0.50, and selection rate 0.25. This represents the number of ways to select 25 people from 100 people, where 50 people are qualified and 21 qualified people are selected. Intuitively, the number of equal-utility allocations is maximized when both: (1) the number of qualified people selected is around half the number of qualified people in the population, and (2) the number of unqualified people selected is around half the number of unqualified people in the population. This follows from the fact that binomial coefficients of the form $n\choose k$ are maximized when $k\approx\frac{n}{2}$.
@@ 4. Why Does Allocation Multiplicity Matter?
Now that we have seen how many equal-utility allocations are possible, we discuss why allocation multiplicity matters. First, risk-averse decision-makers might assume that all or most utility-maximizing allocations can only be produced by top $k$ mappings from near-optimal models. Allocation multiplicity highlights the existence of “riskier” allocations that could have been chosen without sacrificing utility. Second, decision-makers who prioritize normative ideals other than utility may value allocation multiplicity for reasons aligned with their principles. Third, those who view the Rashomon set positively for its potential benefits may find that the space of all equal-utility allocations is what actually delivers those benefits.
4.1. From the Perspective of Normative Ideals
From the Perspective of Risk-Aversion
@@ 4.1. From the Perspective of Normative Ideals
Risk-averse decision-makers will usually choose the allocation with the highest expected utility based on a model’s predictions. From the perspective of the information they have at the time, this might be the most utility-maximizing step they can take. But model multiplicity means that even holding fixed a problem formulation and a training dataset, there are many models that minimize risk to an equivalent extent, thereby delivering equivalent expected utility allocations for the risk averse. However, realized utility, not expected utility, is what ultimately matters to most decision-makers. While it may be difficult to recover all allocations with the same realized utility based on currently available information, the knowledge that additional allocations exist should spur decision-makers to collect the information needed to recover more and sample better from the space of equal-utility allocations. Using a validation dataset to compare the highest-expected-utility allocations with all allocations yielding the same realized utility provides an evaluation framework that may uncover shortcomings of the current training data and inherent limitations of allocating solely on expected risk minimization.
@@
[…]
Evaluating Promises of the Rashomon Set from the Perspective of Allocation Multiplicity: (a) the least discriminatory equal-utility allocation may not be recovered; (b) outcome homogenization may persist in recovered allocations.
From the Perspective of Claims
Allocation multiplicity has a positive role in ensuring that the use of machine learning models to make decisions lives up to our normative commitments. For example, some decision-makers are morally committed to fulfilling the claims (Broome, 1990) that individuals have to goods or outcomes for reasons of fairness. According to the claims view, an individual has a claim to an organ if she needs it more than others; a claim to a job if she merits it; or a claim to a good if she was promised it and therefore deserves it. Claims should be satisfied in proportion to their strength (Broome, 1990). For example, if claims are based on qualification, then all equally qualified individuals should have an equal chance of being selected. Comparing recovered allocations from the Rashomon set to the space of all equal-utility allocations shows that we over-respect the claims of those who appear less risky and under-respect the claims of those who appear riskier. Although we may not always have the information needed to satisfy claims in proportion to their strength, the claims framework suggests that we should (1) seek out additional information that would allow us to better respect all claims, and (2) be prepared to give “compensatory” goods to those whose claims we failed to respect.
From the Perspective of the Worst-Off
Prioritarianism holds that when allocating resources, we should prioritize improvements to the well-being of the worst-off (Arneson, 2022). Within the space of equal-utility allocations, some allocations will give more resources to those who suffer the most. Thus, the prioritarian may want to choose an equal-utility allocation that prioritizes the worst-off, similar to the perspective of choosing the “least discriminatory” algorithm among a set of equally-good models (Black et al., 2024b). However, the shift in perspective from model multiplicity to allocation multiplicity suggests that many of the “multiple” models have allocations that are concentrated on the least-risky individuals. These individuals are likely to be the best-off, not the worst-off. Prioritarianism would suggest going beyond the Rashomon set to find the equal-utility allocations that choose the seemingly-risky people who would be most benefited by receiving positive outcomes.